# algorithmic

### v1

@@ Abstract
The denoising process of diffusion models can be interpreted as a projection of noisy samples onto the data manifold. Moreover, the noise level in these samples approximates their distance to the underlying manifold. Building on this insight, we propose a novel method to enhance sample generation by aligning the estimated noise level with the true distance of noisy samples to the manifold. Specifically, we introduce a noise level correction network, leveraging a pre-trained denoising network, to refine noise level estimates during the denoising process. Additionally, we extend this approach to various image restoration tasks by integrating task-specific constraints, including inpainting, deblurring, super-resolution, colorization, and compressed sensing. Experimental results demonstrate that our method significantly improves sample quality in both unconstrained and constrained generation scenarios. Notably, the proposed noise level correction framework is compatible with existing denoising schedulers (e.g., DDIM), offering additional performance improvements.
@@ 1 Introduction
Generative models have significantly advanced our capability of creating high-fidelity data samples across various domains such as images, audio, and text Song & Ermon (2019); Sohl-Dickstein et al. (2015); Oussidi & Elhassouny (2018). Among these, diffusion models have emerged as one of the most powerful approaches due to their superior performance in generating high-quality samples from complex distributions Song & Ermon (2019); Sohl-Dickstein et al. (2015); Li et al. (2024a). Unlike previous generative models, such as generative adversarial networks (GANs) Goodfellow et al. (2014) and variational autoencoders (VAEs) Kingma & Welling (2014), diffusion models adds multiple levels of noise to the data, and the original data is recovered through a learned denoising process Ho et al. (2020); Song et al. (2023). This allows diffusion models to handle high-dimensional, complex data distributions, making them especially useful for tasks where sample quality and diversity are critical Cao et al. (2024). Their capability to generate complex, high-resolution data has led to widespread applications across numerous tasks, from image generation in models like DALL·E Ramesh et al. (2022) and Stable Diffusion Rombach et al. (2022) to use in robotic path-planningand control Janner et al. (2022); Chi et al. (2023), as well as text generation Li et al. (2022).
@@
The expressive capabilities of diffusion models have also made them a compelling choice for image restoration tasks, where generating high-quality, detailed images is essential Wang et al. (2024). Diffusion models can be used as an image-prior for capturing the underlying structure of image manifold and have shown significant promise for constrained generation such as image restoration Song et al. (2022); Chung et al. (2023a); Wu et al. (2024). Plug-and-play methods were proposed to utilize pre-trained models without the need for extensive retraining or end-to-end optimization for linear inverse problems such as super-resolution, inpainting and compressed sensing Chung et al. (2023a); Dou & Song (2024); Wang et al. (2023). These methods can be interpreted as alternating taking gradient steps towards the constraint set (linear projection for linear inverse problems) and the image manifold (noise direction estimated by the learned denoiser) to find their intersection. However, they may suffer from inconsistency issues if, after each projection step, the noise level no longer approximates distance (Figure 2). Thus correcting the noise level after each step could increase the accuracy of image restoration tasks.
@@
In this work, we propose a novel noise level correction method to refine the estimated noise level and enhance sample generation quality. Our approach introduces a noise level correction network that aligns the estimated noise level of noisy samples more closely with their true distance to the data manifold. By dynamically adjusting the sampling step size based on this corrected noise level estimation, our method improves the sample generation process, significantly enhancing the quality of generated data. Furthermore, our approach integrates seamlessly with existing denoising scheduling methods, such as DDPM (Denoising Diffusion Probabilistic Models) Ho et al. (2020), DDIM (Denoising Diffusion Implicit Models) Song et al. (2021a), and EDM Karras et al. (2022). Furthermore, we extend the application of noise level correction to various image restoration tasks, showing its ability to improve the performance of diffusion-based models such as DDNM (Diffusion Null-Space Model) Wang et al. (2023). Our method achieves improved results across tasks including inpainting, deblurring, super-resolution, colorization, and compressed sensing, as illustrated in fig. 1. Additionally, we introduce a parameter-free lookup table as an approximation of the noise level correction network, providing a computationally efficient alternative for improving the performance of unconstrained diffusion models. In summary, our contributions are:
@@
  • Through extensive experiments, we demonstrate that the proposed noise level correction method consistently provides additional performance gains when applied on top of various denoising methods
@@ 2.1 Diffusion Models
Diffusion models represent a powerful class of latent variable generative models that treat datasets as samples from a probability distribution, typically assumed to lie on a low-dimensional manifold $\mathcal{K}\subset\mathbb{R}^{n}$ Sohl-Dickstein et al. (2015); Ho et al. (2020). Given a data point $z_{0}\in\mathcal{K}$, diffusion models aim to learn a model distribution $p_{\theta}(z_{0})$ that can approximate this manifold and enable the generation of high-quality samples. The process involves gradually corrupting the data with noise during a forward diffusion process and incrementally denoising it to reconstruct the original data through a reverse generative process.
@@
where the noise schedule $\alpha_{t}$ controls the amount of noise injected at each step. Typically, $1\geq\alpha_{1}>\alpha_{2}>\dots>\alpha_{T}\geq 0$, ensuring that $p(z_{T})\sim\mathcal{N}(0,I)$ for large enough $T$ Ho et al. (2020). For mathematical convenience, a reparameterization is often employed, defining new variables $x_{t}=z_{t}/\sqrt{\alpha_{t}}$, which results in:
@@
Where $\sigma_{t}$ denotes the noise level. Note that the diffusion process is originally presented in variable $z_{t}$, we use the formulation $x_{t}=\frac{z_{t}}{\sqrt{\alpha_{t}}}$ to simplify the forward diffusion process and lays the groundwork for the reverse denoising process. A similar formulation can be found in Song et al. (2021b); Karras et al. (2022).
@@
Denoising (sampling) process. The one-step estimation, eq. 5, may lack accuracy, in which case the trained denoiser is applied iteratively through the denoising process. This process aims to progressively denoise a noisy sample $x_{T}$ and recover the original data $x_{0}$. Sampling algorithms construct a sequence of intermediate estimates $(x_{T},x_{T-1},\dots,x_{0})$, starting from an initial point $x_{T}$ drawn from a Gaussian distribution, $x_{T}\sim\mathcal{N}(0,I)$. One of the widely used samplers, the deterministic DDIM Song et al. (2021a), follows the recursion:
@@ 2.2 Additional Related Works
Diffusion models have gained significant attention for their ability to learn complex data distributions, excelling in diverse applications such as image generation Ho et al. (2020); Arechiga et al. (2023), audio synthesis Kong et al. (2021), and robotics Chi et al. (2023). Several methods have been proposed to improve the quality of generated samples while reducing the number of iterations. These include techniques like distillation for reducing sampling steps Nichol & Dhariwal (2021), progressive distillation Salimans & Ho (2022), consistency models Song et al. (2023), and improved design space of diffusion models Karras et al. (2022). On the theoretical side, significant research has explored the non-asymptotic convergence rates of various diffusion samplers, including DDPM Benton et al. (2024) and DDIM Li et al. (2024b), contributing to a deeper understanding of the optimization processes underlying diffusion-based models.
Diffusion models have also demonstrated effectiveness in image restoration tasks, including super-resolution Chung et al. (2023a), inpainting Kawar et al. (2022), deblurring, and compressed sensing Wang et al. (2023). These tasks often require strict adherence to data consistency constraints, making diffusion models particularly suitable for addressing such challenges. Prior works such as DDRM Kawar et al. (2022) and DDNM Wang et al. (2023) pioneered the use of diffusion models for solving linear inverse problems by projecting noisy samples onto the subspace of solutions that satisfy linear constraints. Other methods have employed alternative approaches, such as computing full projections by enforcing linear constraints or using the gradient of quadratically penalized constraints Chung et al. (2022; 2023b). Recent work has extended the application of diffusion models to more complex non-convex constraint functions. In these cases, iterative methods such as gradient descent are utilized to guide the sampling process toward satisfying the constraints, as demonstrated in Universal Guidance Bansal et al. (2023). Furthermore, a provably robust framework for score-based diffusion models applied to image reconstruction was introduced by Xu & Chi (2024), offering robust performance in handling nonlinear inverse problems while ensuring consistency with the observed data.
@@ 3.1 Noise Level as Distance from Noisy Samples to the Manifold
The Manifold Hypothesis suggests that many real-world datasets lie approximately on low-dimensional manifolds embedded in high-dimensional spaces Bengio et al. (2013); Fefferman et al. (2016). In this context, we assume that $\mathcal{K}$ is a manifold of dimension $d$, where $d\ll n$. Leveraging this assumption, we outline the following informal theorem, which summarizes results from Permenter & Yuan (2024) (for formal proofs refer to Permenter & Yuan (2024)):
Theorem 1(Informal) Let $x_{0}\in\mathcal{K}$, $\sigma_{t}>0$, and $\epsilon\sim\mathcal{N}(0,I)$. Define $x_{t}=x_{0}+\sigma_{t}\epsilon\in\mathbb{R}^{n}$. Assuming the manifold hypothesis holds:
(a) Denoising as Approximate Projection:
$\|(x_{t}-\sigma_{t}\epsilon_{\theta}(x_{t},\sigma_{t}))-\text{proj}_{\mathcal{K}}(x_{t})\|\leq\eta\text{dist}_{\mathcal{K}}(x_{t})$ (12)
when $(x_{t},\sigma_{t})$ satisfies: $\frac{1}{\nu}\text{dist}_{\mathcal{K}}(x_{t})\leq\sqrt{n}\sigma_{t}\leq\nu\text{dist}_{\mathcal{K}}(x_{t})$ for constants $1>\eta\geq 0$ and $\nu\geq 1$.
(b) Noise Level as Approximate Distance:
$\sqrt{n}\sigma_{t}\approx\text{dist}_{\mathcal{K}}(x_{t})$ (13)
(c) Projection Interpretation via DDIM: Assume $\epsilon^{\ast}_{\theta}(x,\sigma)$ is the zero-error denoiser, and the initial distance satisfies $\text{dist}_{\mathcal{K}}(x_{T})=\sqrt{n}\sigma_{T}$. Then the DDIM sampler generates the sequence ($x_{T},\dots,x_{0}$) by performing gradient descent on the objective function $f(x):=\frac{1}{2}\text{dist}_{\mathcal{K}}(x)^{2}$ with a step-size of $\beta_{t}:=1-\sigma_{t-1}/\sigma_{t}$:
@@
Theorem 1(a) demonstrates that the estimated clean sample generated by the denoising process $\hat{x}_{0|t}=x_{t}\sigma_{t}\epsilon_{\theta}(x_{t},\sigma_{t}))$, serves as an approximation of the projection of the noisy sample $x_{t}$ onto the manifold $\mathcal{K}$. Similarly, Theorem 1(b) establishes that the distance of a noisy sample $x_{t}$ onto manifold $\mathcal{K}$, can be approximated by the noise level $\sqrt{n}\sigma_{t}$. With a zero-error denoiser, Theorem 1(c) shows that the DDIM process generates a sequence that monotonically decreases the distance to the manifold by following the gradient descent of the objective function $f(x)$. Specifically, the distance of each noisy sample from the manifold is determined by the noise level, $\text{dist}_{\mathcal{K}}(x_{t})=\sqrt{n}\sigma_{t}$, while the direction of the denoising process, or projection, is guided by the estimated noise vector $\nabla\text{dist}_{\mathcal{K}}(x_{t})=\epsilon^{\ast}_{\theta}(x_{t},\sigma_{t})/\sqrt{n}$.
@@
However Theorem 1(c), eq. 15, relies on two key assumptions: (1) an zero-error denoiser, where $\epsilon_{\theta}(x_{t},\sigma_{t})=\epsilon_{t}$, and (2) an initial condition where the distance at $t=T$, $\text{dist}_{\mathcal{K}}(x_{T})=\sqrt{n}\sigma_{T}$. For sample generation, cumulative errors introduced during the denoising process can lead to biases due to the imperfections of the denoiser Ning et al. (2023). These errors become particularly significant by the final steps, where deviations in the denoising process result in a large mismatch between the true distance $\text{dist}_{\mathcal{K}}(x_{t})$ and the estimated noise level $\sqrt{n}\sigma_{t}$, as illustrated in Figure 2a. As observed, a mismatch between the true distance and the estimated noise level can lead to the clean image estimation using eq. 5 falling outside the manifold $\mathcal{K}$. In constrained sample generation tasks, such as image restoration and guided sample generation, this issue is further exacerbated. As shown in Figure 2b, deviations introduced by guidance terms or projection steps onto the constraint $\mathcal{C}$ amplify the discrepancy between the estimated noise level $\sqrt{n}\sigma_{t}$ and the actual distance $\text{dist}_{\mathcal{K}}(x_{t})$. These deviations are particularly impactful in the later denoising steps and can prevent the reconstructed sample from accurately lying on the data manifold $\mathcal{K}$. Additional details are provided in Appendix A.

### v2

@@ Abstract
The denoising process of diffusion models can be interpreted as an approximate projection of noisy samples onto the data manifold. Moreover, the noise level in these samples approximates their distance to the underlying manifold. Building on this insight, we propose a novel method to enhance sample generation by aligning the estimated noise level with the true distance of noisy samples to the manifold. Specifically, we introduce a noise level correction network, leveraging a pre-trained denoising network, to refine noise level estimates during the denoising process. Additionally, we extend this approach to various image restoration tasks by integrating task-specific constraints, including inpainting, deblurring, super-resolution, colorization, and compressed sensing. Experimental results demonstrate that our method significantly improves sample quality in both unconstrained and constrained generation scenarios. Notably, the proposed noise level correction framework is compatible with existing denoising schedulers (e.g., DDIM), offering additional performance improvements.
@@ 1 Introduction
Generative models have significantly advanced our capability of creating high-fidelity data samples across various domains such as images, audio, and text (Song & Ermon, 2019; Sohl-Dickstein et al., 2015; Oussidi & Elhassouny, 2018). Among these, diffusion models have emerged as one of the most powerful approaches due to their superior performance in generating high-quality samples from complex distributions (Song & Ermon, 2019; Sohl-Dickstein et al., 2015; Li et al., 2024a). Unlike previous generative models, such as generative adversarial networks (GANs) (Goodfellow et al., 2014) and variational autoencoders (VAEs) (Kingma & Welling, 2014), diffusion models add multiple levels of noise to the data, and the original data is recovered through a learned denoising process (Ho et al., 2020; Song et al., 2023). This allows diffusion models to handle high-dimensional, complex data distributions, making them especially useful for tasks where sample quality and diversity are critical (Cao et al., 2024). Their capability to generate complex, high-resolution data has led to widespread applications across numerous tasks, from image generation in models like DALL·E (Ramesh et al., 2022) and Stable Diffusion (Rombach et al., 2022) to use in robotic path-planningand control (Janner et al., 2022; Chi et al., 2023), as well as text generation (Li et al., 2022).
@@
The expressive capabilities of diffusion models have also made them a compelling choice for image restoration tasks, where generating high-quality, detailed images is essential (Wang et al., 2024). Diffusion models can be used as an image-prior for capturing the underlying structure of image manifold and have shown significant promise for constrained generation such as image restoration (Song et al., 2022; Chung et al., 2023a; Wu et al., 2024). Plug-and-play methods were proposed to utilize pre-trained models without the need for extensive retraining or end-to-end optimization for linear inverse problems such as super-resolution, inpainting and compressed sensing (Chung et al., 2023a; Dou & Song, 2024; Wang et al., 2023). These methods can be interpreted as alternating taking gradient steps towards the constraint set (linear projection for linear inverse problems) and the image manifold (noise direction estimated by the learned denoiser) to find their intersection. However, they may suffer from inconsistency issues if, after each projection step, the noise level no longer approximates distance (Figure 2). Thus correcting the noise level after each step could increase the accuracy of image restoration tasks.
@@
In this work, we propose a novel noise level correction method to refine the estimated noise level and enhance sample generation quality. Our approach introduces a noise level correction network that aligns the estimated noise level of noisy samples more closely with their true distance to the data manifold. By dynamically adjusting the sampling step size based on this corrected noise level estimation, our method improves the sample generation process, significantly enhancing the quality of generated data. Furthermore, our approach integrates seamlessly with existing denoising scheduling methods, such as DDPM (Denoising Diffusion Probabilistic Models) (Ho et al., 2020), DDIM (Denoising Diffusion Implicit Models) (Song et al., 2021a), and EDM (Karras et al., 2022). Furthermore, we extend the application of noise level correction to various image restoration tasks, showing its ability to improve the performance of diffusion-based models such as DDNM (Denoising Diffusion Nullspace Model) in Wang et al. (2023). Our method achieves improved results across tasks including inpainting, deblurring, super-resolution, colorization, and compressed sensing, as illustrated in Figure 1. Additionally, we introduce a parameter-free lookup table as an approximation of the noise level correction network, providing a computationally efficient alternative for improving the performance of unconstrained diffusion models. In summary, our contributions are:
@@
  • Through extensive experiments, we demonstrate that the proposed noise level correction method consistently provides additional performance gains when applied on top of various denoising methods.
@@ 2.1 Diffusion Models
Diffusion models represent a powerful class of latent variable generative models that treat datasets as samples from a probability distribution, typically assumed to lie on a low-dimensional manifold $\mathcal{K}\subset\mathbb{R}^{n}$ (Sohl-Dickstein et al., 2015; Ho et al., 2020). Given a data point $z_{0}\in\mathcal{K}$, diffusion models aim to learn a model distribution $p_{\theta}(z_{0})$ that can approximate this manifold and enable the generation of high-quality samples. The process involves gradually corrupting the data with noise during a forward diffusion process and incrementally denoising it to reconstruct the original data through a reverse generative process.
@@
where the noise schedule $\alpha_{t}$ controls the amount of noise injected at each step. Typically, $01>\alpha_{1}>\alpha_{2}>\dots>\alpha_{T}>$ 0, ensuring that $p(z_{T})\sim\mathcal{N}(0,I)$ for large enough $T$ (Ho et al., 2020). For mathematical convenience, a reparameterization is often employed, defining new variables $x_{t}=z_{t}/\sqrt{\alpha_{t}}$, which results in:
@@
Where $\sigma_{t}$ denotes the noise level. Note that the diffusion process is originally presented in variable $z_{t}$, we use the formulation $x_{t}=\frac{z_{t}}{\sqrt{\alpha_{t}}}$ to simplify the forward and reverse denoising processes. A similar formulation can be found in Song et al. (2021b); Karras et al. (2022).
@@
Denoising (sampling) process. The one-step estimation, eq. 5, may lack accuracy, in which case the trained denoiser is applied iteratively through the denoising process. This process aims to progressively denoise a noisy sample $x_{T}$ and recover the original data $x_{0}$. Sampling algorithms construct a sequence of intermediate estimates $(x_{T},x_{T-1},\dots,x_{0})$, starting from an initial point $x_{T}$ drawn from a Gaussian distribution, $x_{T}\sim\mathcal{N}(0,I)$. One of the widely used samplers, the deterministic DDIM (Song et al., 2021a), follows the recursion:
@@ 2.2 Additional Related Works
Diffusion models have gained significant attention for their ability to learn complex data distributions, excelling in diverse applications such as image generation (Ho et al., 2020; Arechiga et al., 2023), audio synthesis (Kong et al., 2021), and robotics (Chi et al., 2023). Several methods have been proposed to improve the quality of generated samples while reducing the number of iterations. These include techniques such as distillation to reduce sampling steps (Nichol & Dhariwal, 2021), progressive distillation (Salimans & Ho, 2022), consistency models (Song et al., 2023), and improved design space of diffusion models (Karras et al., 2022). On the theoretical side, significant research has explored the non-asymptotic convergence rates of various diffusion samplers, including DDPM (Benton et al., 2024) and DDIM (Li et al., 2024b), contributing to a deeper understanding of the optimization processes underlying diffusion-based models.
Diffusion models have also demonstrated effectiveness in image restoration tasks, including super-resolution (Chung et al., 2023a), inpainting (Kawar et al., 2022), deblurring, and compressed sensing (Wang et al., 2023). These tasks often require strict adherence to data consistency constraints, making diffusion models particularly suitable for addressing such challenges. Prior works such as DDRM (Kawar et al., 2022) and DDNM (Wang et al., 2023) pioneered the use of diffusion models to solve linear inverse problems by projecting noisy samples onto the subspace of solutions that satisfy linear constraints. Other methods have employed alternative approaches, such as computing full projections by enforcing linear constraints or using the gradient of quadratically penalized constraints (Chung et al., 2022; 2023b). Recent work has extended the application of diffusion models to more complex non-convex constraint functions. In these cases, iterative methods such as gradient descent are utilized to guide the sampling process toward satisfying the constraints, as demonstrated in Universal Guidance (Bansal et al., 2023). Furthermore, a provably robust framework for score-based diffusion models applied to image reconstruction was introduced by Xu & Chi (2024), offering robust performance in handling nonlinear inverse problems while ensuring consistency with the observed data.
@@ 3.1 Noise Level as Distance from Noisy Samples to the Manifold
Permenter & Yuan (2024) introduces the following relative-error model, inspired by the manifold hypothesis (Bengio et al., 2013; Fefferman et al., 2016) as well as from the fact that the ideal/optimal denoiser is the gradient of a smoothed distance function:
Assumption 3.1 (Informal)
The learned denoiser $\epsilon^{*}$ as well as the sampling noise schedule $\{\sigma_{t}\}_{t=1}^{N}$ satisfies the following for all $x_{t}$ during sampling:
  1. 1.
    $\sigma_{t}$ decreases slowly enough to ensure that $\sqrt{n}\sigma_{t}\approx\text{dist}_{\mathcal{K}}(x_{t})$ after each iteration
  2. 2.
    If $\sigma_{t}$ approximates $\text{dist}_{\mathcal{K}}(x_{t})$, then $x_{t}-\sigma_{t}\epsilon^{*}(x_{t},\sigma_{t})\approx\text{proj}_{\mathcal{K}}(x_{t})$ with constant relative error with respect to $\text{dist}_{\mathcal{K}}(x_{t})$.
Leveraging this assumption, we outline the following theorem informally, which summarizes results from Permenter & Yuan (2024) (for formal proofs refer to Permenter & Yuan (2024)):
Theorem 3.1 (Informal)
Suppose Assumption 3.1 holds and the initial distance satisfies $\text{dist}_{\mathcal{K}}(x_{T})=\sqrt{n}\sigma_{T}$. Then the DDIM sampler generates the sequence ($x_{T},\dots,x_{0}$) by performing gradient descent on the objective function $f(x):=\frac{1}{2}\text{dist}_{\mathcal{K}}(x)^{2}$ with a step-size of $\beta_{t}:=1-\sigma_{t-1}/\sigma_{t}$:
@@
Theorem 3.1 demonstrates that the estimated clean sample generated by the denoising process $\hat{x}_{0|t}=x_{t}\sigma_{t}\epsilon_{\theta}(x_{t},\sigma_{t}))$ serves as an approximation of the projection of the noisy sample $x_{t}$ onto the manifold $\mathcal{K}$. It also establishes that throughout the sampling process, the distance of a noisy sample $x_{t}$ onto manifold $\mathcal{K}$, can be approximated by the noise level $\sqrt{n}\sigma_{t}$. Given that $\sigma_{t}$ and $\epsilon_{\theta}$ satisfies Assumption 3.1, the DDIM process generates a sequence $d_{t}$ with monotonically decreasing distance to the manifold by gradient descent on the squared-distance function to the manifold. Specifically, the distance of each noisy sample from the manifold is determined by the noise level, $\text{dist}_{\mathcal{K}}(x_{t})=\sqrt{n}\sigma_{t}$, while the gradient $\nabla f(x)$, which is the direction of projection of $x_{t}$ onto the manifold, is estimated by the noise vector $\nabla\text{dist}_{\mathcal{K}}(x_{t})\approx\epsilon^{\ast}_{\theta}(x_{t},\sigma_{t})/\sqrt{n}$.
@@
However Theorem 3.1 relies on two key assumptions: (1) a denoiser $\epsilon_{\theta}(x_{t},\sigma_{t})$ satisfying the relative projection-error assumption, and (2) a noise schedule that decrease slowly enough so that for all $t$, we have $\text{dist}_{\mathcal{K}}(x_{t})\approx\sqrt{n}\sigma_{t}$. For sample generation, cumulative errors introduced during the denoising process can lead to biases due to the imperfections of the denoiser Ning et al. (2023). These errors become particularly significant by the final steps, where deviations in the denoising process result in a large mismatch between the true distance $\text{dist}_{\mathcal{K}}(x_{t})$ and the estimated noise level $\sqrt{n}\sigma_{t}$, as illustrated in Figure 2a. As observed, a mismatch between the true distance and the estimated noise level can lead to the clean image estimation using eq. 5 falling outside the manifold $\mathcal{K}$. In constrained sample generation tasks, such as image restoration and guided sample generation, this issue is further exacerbated. As shown in Figure 2b, deviations introduced by guidance terms or projection steps onto the constraint $\mathcal{C}$ amplify the discrepancy between the estimated noise level $\sqrt{n}\sigma_{t}$ and the actual distance $\text{dist}_{\mathcal{K}}(x_{t})$. These deviations are particularly impactful in the later denoising steps and can prevent the reconstructed sample from accurately lying on the data manifold $\mathcal{K}$. Additional details are provided in Appendix A.